OpenStudy (anonymous):

According to The Fundamental Theorem of Algebra, how many zeros does the function f(x) = 15x23 + 41x19 + 13x5– 10 have

jimthompson5910 (jim_thompson5910):

The fundamental theorem of algebra states that a polynomial function of degree n will have n complex zeros The degree of f(x) = 15x^23 + 41x^19 + 13x^5– 10 is 23 since that is the largest exponent So there are 23 complex zeros (or just 23 zeros in general)

jimthompson5910 (jim_thompson5910):

There's not much to do here but just memorize what the theorem is saying

OpenStudy (anonymous):

ok lol i really hate mathhh!

jimthompson5910 (jim_thompson5910):

well with more practice, it will get easier

OpenStudy (anonymous):

haha if you say so... lets try another one okay What are the zeros of the polynomial function: f(x) = x3 + 2x2 – 24x ?

jimthompson5910 (jim_thompson5910):

Can you factor that at all?

OpenStudy (anonymous):

do i just combine like terms or something.

jimthompson5910 (jim_thompson5910):

What do x^3, 2x^2 and -24x all have in common?

OpenStudy (anonymous):

they all have x?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so we can factor this out f(x) = x^3 + 2x^2 – 24x f(x) = x(x^2 + 2x – 24)

jimthompson5910 (jim_thompson5910):

Now find two numbers that both multiply to -24 AND add to 2 at the same time

OpenStudy (anonymous):

-12 and 2 ?

jimthompson5910 (jim_thompson5910):

they multiply to -24, but do NOT add to 2

jimthompson5910 (jim_thompson5910):

what else

OpenStudy (anonymous):

ohhh wow my bad haha hold on ...

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

-6 and 4

jimthompson5910 (jim_thompson5910):

closer

jimthompson5910 (jim_thompson5910):

but notice how -6+4 = -2

OpenStudy (anonymous):

-4 and 6 ((:

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

so x^2+2x-24 factors to (x-4)(x+6)

OpenStudy (anonymous):

so i could have just done the diamond problem? haha

jimthompson5910 (jim_thompson5910):

so we can say f(x) = x^3 + 2x^2 – 24x f(x) = x(x^2 + 2x – 24) f(x) = x(x-4)(x+6) 0 = x(x-4)(x+6) x(x-4)(x+6) = 0 x = 0, x-4=0 or x+6 = 0

jimthompson5910 (jim_thompson5910):

yes you could factored using the diamond method

jimthompson5910 (jim_thompson5910):

could have factored*

OpenStudy (anonymous):

okayyyyyy....lol

jimthompson5910 (jim_thompson5910):

now solve for x in each equation

OpenStudy (anonymous):

6,0,-4

jimthompson5910 (jim_thompson5910):

oh wait, i went too fast lol

jimthompson5910 (jim_thompson5910):

f(x) = x^3 + 2x^2 – 24x f(x) = x(x^2 + 2x – 24) f(x) = x(x-4)(x+6) 0 = x(x-4)(x+6) x(x-4)(x+6) = 0 x = 0, x-4=0 or x+6 = 0 x = 0, x-4+4=0+4 or x+6-6 = 0-6 x = 0, x=4 or x = -6 So the zeros of f(x) = x^3 + 2x^2 – 24x are: -6,0,4

OpenStudy (anonymous):

yayyyyyyy i did it (((:

jimthompson5910 (jim_thompson5910):

you were close, but it should be -6 instead of +6

jimthompson5910 (jim_thompson5910):

so you got what I got?

OpenStudy (anonymous):

yesss

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (anonymous):

What are the possible rational zeros of f(x) = x4 + 2x3 – 3x2 – 4x + 20 ? what on earthhhh? i had subs my whole junior year in math):

jimthompson5910 (jim_thompson5910):

List all the possible factors of 20

jimthompson5910 (jim_thompson5910):

What are they

OpenStudy (anonymous):

UHHHHHH....

OpenStudy (anonymous):

2 AND 10

jimthompson5910 (jim_thompson5910):

good, what else

jimthompson5910 (jim_thompson5910):

20 = 2*10

OpenStudy (anonymous):

5 and 4

jimthompson5910 (jim_thompson5910):

good 20 = 5*4

jimthompson5910 (jim_thompson5910):

anything else?

OpenStudy (anonymous):

20 and 1

jimthompson5910 (jim_thompson5910):

20 = 1*20

jimthompson5910 (jim_thompson5910):

so we have 1*20, 2*10 4*5

jimthompson5910 (jim_thompson5910):

Therefore, the factors of 20 are 20: 1, 2, 4, 5, 10, 20

jimthompson5910 (jim_thompson5910):

Now divide each of these factors by the factors of the leading coefficient (which in this case, it's just 1) 1/1 = 1 2/1 = 2 4/1 = 4 5/1 = 5 10/1 = 10 20/1 = 20 So we get the list 1, 2, 4, 5, 10, 20 Now add on the negative of each number listed 1, 2, 4, 5, 10, 20 -1, -2, -4, -5, -10, -20

jimthompson5910 (jim_thompson5910):

So all the possible rational roots are 1, 2, 4, 5, 10, 20 -1, -2, -4, -5, -10, -20

OpenStudy (anonymous):

ohhh okay wow these seem hard but there really not if you know how to start them.

jimthompson5910 (jim_thompson5910):

exactly, starting is the hardest part usually (for me anyways)

OpenStudy (anonymous):

hahaha mee toooo!

jimthompson5910 (jim_thompson5910):

that's why it helps to start at the end sometimes (ie start with what you want to get and see how to get there)

jimthompson5910 (jim_thompson5910):

it's a bit odd, but it works sometimes

OpenStudy (anonymous):

ive never thought about that. lol

jimthompson5910 (jim_thompson5910):

it's a very useful trick that took me a while to grasp lol (i didn't learn it til recently either...and i still have trouble grasping it sometimes)